# An Evaluation of Borda Count Variations Using Ranked Choice Voting Data

### v1

@@ 1 Introduction
The vast majority of elections in the United States have been conducted using the single-vote plurality method, other than the use of the Electoral College for Presidential elections. Increasingly over the past twenty years, ranked choice voting (RCV), particularly through the method of instant runoff voting (IRV), has been implemented across numerous municipal and statewide elections including Presidential primaries and U.S. Representative elections in states such as Alaska and Maine. Both of these methods are vulnerable to multiple voting failures. plurality is especially problematic, since ranked ballots are necessary to determine if some failures, such as the majority loser or Condorcet failures have occurred. The following are voting failures, which will be defined in the next section, that can occur with these two methods:
@@
  • RCV can exhibit upward and downward monotonicity paradoxes.
  • RCV can have a truncation failure through occurrences of the no-show paradox.
Recent work has focused on comparing results of various voting methods using U.S. elections conducted with IRV. McCune and McCune (2024) showed that each pair of methods agreed on a winning candidate in the vast majority of cases with the lowest pair being the plurality method compared to a variation of the Borda count which produced the same winner about $76\%$ of the time. Graham-Squire and McCune (2022) examined a similar set of RCV elections that used the IRV method to determine the rates at which voting failures occur, finding that while uncommon, IRV elections sometimes exhibit the various failures listed above. Arrow (1963) famously proved in his Impossibility Theorem that no voting method exists that can always avoid these failures, but we can still strive to limit their occurrences and particularly eliminate the most undesirable paradoxes.
@@ 2 Voting Failures
We begin by defining the primary means of evaluating the fairness of election methods, which is through the observation of voting failures, sometimes referred to in the positive sense as fairness criteria. We know from Arrow’s Impossibility Theorem that every voting method can exhibit at least one of a small subset of these failures. While a perfect method that avoids every failure does not exist, we will demonstrate which failures are possible or will never occur when using the Borda count variations, which will be aided by grouping the failures into three distinct sets.
@@
    • Majority Winner/Loser: The majority winner failure occurs when a candidate has a majority of first-place votes but loses the election. The majority loser failure, on the other hand, occurs when a candidate has a majority of last-place votes yet wins the election. Note that partial ballots with more than one candidate unranked will not be considered as having a last-place vote, but are still considered in the total when determining if a majority exists.
    • Condorcet Winner/Loser: The Condorcet winner failure occurs if a candidate would defeat every other candidate in a pairwise comparison, but loses the election. The Condorcet loser failure occurs if a candidate loses to every other candidate in a head-to-head comparison, yet wins the election. With partial ballots, note that a candidate ranked on the partial ballot is clearly favored over all unranked candidates, but partial ballots are not considered in a pairwise comparison if both candidates are unranked on that ballot.
@@
    • No-Show Paradox: The no-show paradox is an extreme form of the soon-to-be defined truncation failure that occurs if there a set a voters whose top preference is a losing candidate, but removing those voters would result in that candidate now winning the election. Hence those voters have a more desirable outcome occur by staying home and not voting.
@@ 3 Borda Count Variations
We will examine the outcomes and fairness of several variations of the Borda count method in our upcoming sections, which we will refer to as BC variations. We must first distinguish between two implementations of the standard Borda count points system described by Emerson (2013) and used by McCune and McCune (2024) to account for partial or truncated ballots. The implementation of the points system introduced by Borda with the points vector $(n,n-1,\ldots,2,1)$ is assuming every ballot is complete; that is, each voter ranked a full list of candidates. In most ranked choice elections with $n$ candidates, however, voters may choose to submit a partial ballot in which they rank fewer than $n$ options. Additionally, some districts limit voters to ranking only a subset of candidates of a maximum size $m<n$. These are called truncated ballots. For example, the Minneapolis City Council elections only allow at most 3 candidates to be included in a ballot no matter how many candidates are running.
@@
The two primary approaches for handling partial or truncated ballots are the Borda count untreated (BCU) and the modified Borda count (MBC). If a ballot only ranks $k$ of the $n$ candidates, MBC awards $(k,k-1,\ldots,2,1,0,\ldots,0)$, so all unranked candidates receive 0 points. BCU awards points to the ranked candidates based on the full Borda count value as $(n,n-1,\ldots,n-k+1,0,\ldots,0)$. Emerson (2013) argues that BCU would encourage voters to use a form of tactical voting in which voters cast a ranked ballot consisting of only their top choice, which is known as bullet voting. Giving a full $n$ points to that choice and 0 points to all other candidates would clearly give a voter’s top candidate the best chances of winning, and if enough voters follow this strategy, the election would devolve into a plurality vote and cause frequent truncation failures.
@@
One negative of both variations is that votes are no longer considered equal. For example, in an election with $5$ candidates, a voter who fills out a complete ballot ranking all 5 candidates is assigning 15 total points to the candidates. A voter with a partial ballot of 3 candidates would only give candidates a total of 12 points using BCU and merely 6 points using MBC. While this does incentivize the electorate to fill out complete ballots, it is against democratic principles to count some voters more than others. Furthermore, in elections such as the 2021 East Hampton Mayor election, as many as 66 candidates may be present on the ballot. In elections with more than 8 candidates, which account for over 10 percent of the RCV elections in our dataset, it would not feasible to expect each voter to differentiate every candidate enough to provide a complete ranking without arbitrarily ordering the lowest set of candidates. MBC in particular penalizes voters who want to submit a partial ballot to the point that they may be pressured to vote insincerely to make sure their top candidate is awarded the maximum number of points.
@@
To address this inequity, we favor the variation of the Borda count introduced by Dummett (1997) called the averaged Borda count (ABC). In the previous example with 5 candidates and a 3-candidate partial ballot, it initially follows BCU with 5 points for 1st, 4 for 2nd, and 3 for 3rd, but ABC will then average the unassigned 3 points to give 1.5 points to each of the unlisted candidates. In general, the $k$ candidates who are ranked in an $n$-candidate election receive the same points as in BCU, but the unranked $n-k$ candidates receive an equal share of the points that a complete ballot would have assigned. This equates to $\frac{1}{n-k}(1+2+\cdots+n-k)=\frac{(n-k)(n-k+1)}{2(n-k)}=\frac{n-k+1}{2}$. In total, the points vector is
@@
The primary two points-scale variations that we will investigate are the Exponential Borda Count (EBC) and Quadratic Borda Count (QBC). EBC awards 1 point to the last ranked candidate and double the points for each subsequently higher rank. This leads to the points for the $i$th candidate being the exponential function $p_{i}=2^{n-i}$, or points vector of $(2^{n-1},2^{n-2},\ldots,2,1)$. QBC also awards 1 point for a last-place vote and then the difference in points between ranks increases by 1 each time with the top two candidates having a difference of $p_{1}-p_{2}=n-1$. In general, $p_{i}-p_{i+1}=p_{i+1}-p_{i+2}+1$, and the formula for the points assigned to the $i$th ranked candidate is the quadratic function
@@
Table 2 shows the values of $p_{i}$ in a 6-candidate election to illustrate the differences in the variations. Observe the differences between pairs of entries in the QBC column increase by 1 as you move up the column from 1 in the last two ranks to 5 between the first two ranks.
@@
Now that we have defined each variation, we shift our focus to whether each voting failure can occur with each method. Table 3 displays this information for the currently used plurality and IRV methods, along with the five Borda variations we are investigating. Any failure that is possibly exhibited by these variations will either occur within the empirical analysis in an upcoming section, or is demonstrated by the referenced example. The failures for which a Borda variation is immune will be proven so in the upcoming theorems, with the exception of MBC avoiding truncation failures, as this was previously shown (Kamwa 2022).
@@
The following are elections of very small voter size that demonstrate particular voting failures for some of our Borda variations. Although these are toy examples to make the points calculation and pairwise comparisons easy, examples can be developed for much larger elections by keeping similar distributions in ballot types. As we will see in Section 6, none of these voting failure types would have occurred using the specified Borda variation on previous U.S. elections, hence the need for examples to demonstrate their possibility. We begin by demonstrating that EBC and QBC can exhibit a majority winner failure.
@@
Note the previous example is also a Condorcet winner failure, but we only focused on the majority aspect because it does not occur in our upcoming empirical analysis, whereas Condorcet winner failures are exhibited for every variation of the Borda count. The next example shows that EBC and QBC can both have majority loser and Condorcet loser failures. Although the Borda count cannot exhibit these two failures, as shown by Fishburn and Gehrlein (1976), any extension of Borda with a points vector where the difference $p_{i}-p_{i-1}$ increases as $i$ goes from $2$ to $n$ can have these failures if the number of voters is sufficiently large.
@@
Consider an election with candidates $A$, $B$, and $C$ using EBC or the equivalent QBC. Assume there are 4 ballots ranking $A>B>C$, 3 ranking $A>C>B$, 4 with $B>C>A$, and 4 with $C>B>A$. Notice that $A$ has last-place votes in 8 of the 15 ballots, which is a majority and leads $A$ to lose in pairwise competition to both $B$ and $C$. Candidate $A$ wins the election, however, because their total points are 36 while $B$ earns 35 and $C$ receives 34. This election exhibits both a majority loser and Condorcet loser failure.
@@
It was shown (Fishburn and Gehrlein 1976) that the Borda count, assuming all ballots are complete, cannot have a Condorcet loser failure. This implies a majority loser failure is also impossible since a majority loser candidate by default will lose all of its pairwise competitions. We now show two variations of Borda, particularly MBC and ABC, are also immune to majority loser failures, and in the case of ABC, Condorcet loser failures as well.
@@
We follow a similar approach to the previous result by supposing for a contradiction that there is an ABC election with a Condorcet loser failure. Assume $A$ is the winning candidate who loses in pairwise competition to all other candidates. The standard Borda count does not have Condorcet loser failures (Fishburn and Gehrlein 1976), so there must exist partial ballots.
@@
In the new Borda count election, all previously complete ballots are included twice. Any ballot where one of $A$ or $B$ was ranked and the other unranked would also count twice toward the pairwise comparison. This is because the originally ranked candidate is favored in both new ballots, despite the order of the newly ranked candidates being different in the two forms of completing ballots. This accounts for all cases where one of $A$ or $B$ was ranked higher, and thus there are so far $2x$ ballots with $B$ favored over $A$, and $2y$ with the opposite relative order. Assuming there were $z$ partial ballots that left $A$ and $B$ unranked, the two forms of ballot completion with reverse ordering result in $A$ ranked higher in $z$ newly completed ballots and $B$ also being higher in $z$ ballots. Now that we have covered all cases of both being ranked, one ranked and one unranked, and both unranked, $B$ is favored over $A$ is $2x+z$ ballots, while $A$ is favored in $2y+z$ ballots. The fact that $x>y$ directly implies $2x+z>2y+z$, meaning that $B$ wins the head-to-head comparison.
@@ 4 Data and Methodology
This election data used in our empirical analysis consists of 421 RCV elections from a publicly available database (Otis 2022). Table 4 summarizes the dataset, including the states and cities where the elections occurred and the range of years for each jurisdiction. The elections include local city council, school board, and mayor offices in addition to statewide elections for the U.S. House of Representatives and Presidential primaries.
In using the RCV data where voters chose their ranking based on IRV deciding the winner, we recognize a potential methodological flaw by which some voter preferences may differ a points-based method was used. Strategic voting, such as truncating a ballot to remove your favored candidate’s top competitor, would nevertheless have a negligible effect on potential verifiable failures. Meanwhile, the purpose of voter manipulation failures is to pinpoint this potential behavior and identify it as an unfair outcome, erasing any methodological concerns.
There are multiple considerations needed for processing ballots and implementing our points-based methods. For instance, there are many ballots in which a candidate was listed multiple times, ranks were skipped, or multiple candidates were listed in the same ranking, identified as an overvote. Any appearance of a candidate after the first instance was removed from the ballot, likewise for skipped or undeclared ranks. These were then processed as partial ballots. Finally, for cases of overvotes, there was no way to determine which candidates were in the overvote, so all candidates below that ranking were removed, leaving only the candidates ranked above the overvote rank.
Another determination that must be made when implementing a points-based voting system is how to enumerate ballots that include write-in candidates. Across all elections in which write-in candidates were allowed to be included, they received no more than $1.6\%$ of the total first-place votes. Moreover, this number includes all write-in candidates collectively since data from elections offices usually reports these as a single candidate named Write-In or a similar designation. Therefore, we opted for removing all write-ins since they had a negligible effect on the election outcome.
@@
The presence of voter/candidate manipulation failures are more difficult to identify from a computational standpoint. For a truncation failure, our algorithm considers each losing candidate and truncates a set of ballots to give them an optimal chance of overtaking the winning candidate’s point total. This is accomplished by removing all ranks starting with the initial winner from each ballot where the particular losing candidate was ranked higher. If that candidate now wins, there is a truncation failure. There are scenarios, however, where a third candidate would now win, forcing us to further analyze those examples to determine if a smaller subset of ballots or if a different truncation strategy, such as removing this third candidate as well, would trigger a failure. Compromise failures follow a similar algorithmic process where a losing candidate ranked higher than the winner is moved to the top of all such ballots to determine if they would now win.
To identify spoiler effects, we use a brute force approach for smaller elections with slates of 10 or fewer candidates. By sequentially removing each subset of losing candidates, the algorithm determines if the winner has changed, resulting in a spoiler effect. Unfortunately, it becomes computationally infeasible to perform this evaluation for every subset for larger slates. To guarantee accuracy, we limited our analysis to the 396 elections with 10 or fewer candidates.

### v2

@@ 1 Introduction
The vast majority of elections in the United States have been conducted using the single-vote plurality method, other than the use of the Electoral College for Presidential elections. Increasingly over the past twenty years, ranked choice voting (RCV), particularly through the method of instant runoff voting (IRV), has been implemented across numerous municipal and statewide elections including Presidential primaries and U.S. Representative elections in states such as Alaska and Maine. Both of these methods are vulnerable to multiple voting failures. Plurality is especially problematic, since ranked ballots are necessary to determine if some failures, such as the majority loser or Condorcet failures, have occurred. The following are voting failures, which will be defined in the next section, that can occur with these two methods:
@@
  • IRV can exhibit upward and downward monotonicity paradoxes.
  • IRV can have a truncation failure through occurrences of the no-show paradox.
Recent work has focused on comparing results of various voting methods using U.S. elections conducted with IRV. McCune and McCune (2024) showed that each pair of methods agreed on a winning candidate in the vast majority of cases with the lowest pair being the plurality method compared to a variation of the Borda count which produced the same winner about $76\%$ of the time. Graham-Squire and McCune (2022) examined a similar set of RCV elections that used the IRV method to determine the rates at which voting failures occur, finding that while uncommon, IRV elections sometimes exhibit the various failures listed above. Arrow (1963) famously proved in his Impossibility Theorem that no voting method exists that will always avoid these failures, but we can still strive to limit their occurrences and particularly eliminate the most undesirable paradoxes.
@@ 2 Voting Failures
We begin by defining the primary means of evaluating the fairness of election methods, which is through the observation of voting failures, sometimes referred to in the positive sense as fairness criteria. It is known from Arrow’s Impossibility Theorem that every voting method can exhibit at least one of a small subset of these failures. We will demonstrate which failures are possible or will never occur when using the Borda count variations, which will be aided by grouping the failures into three distinct sets.
@@
    • Majority Winner/Loser Failure: A majority winner failure occurs when a candidate has a majority of first-place votes but loses the election. A majority loser failure, on the other hand, occurs when a candidate has a majority of last-place votes yet wins the election. Note that partial ballots with more than one candidate unranked will not be considered as having a last-place vote, but are still considered in the total when determining if a majority exists.
    • Condorcet Winner/Loser Failure: A Condorcet winner failure occurs if a candidate would defeat every other candidate in a pairwise comparison, but loses the election. A Condorcet loser failure occurs if a candidate loses to every other candidate in a head-to-head comparison, yet wins the election. With partial ballots, note that a candidate ranked on the partial ballot is clearly favored over all unranked candidates, but partial ballots are not considered in a pairwise comparison if both candidates are unranked on that ballot.
@@
    • No-Show Paradox: The no-show paradox is an extreme form of the soon-to-be defined truncation failure that occurs if there is a set of voters whose top preference is a losing candidate, but removing those ballots would result in that candidate now winning the election. Hence those voters have a more desirable outcome occur by staying home and not voting.
@@ 3 Borda Count Variations
We will examine the outcomes and fairness of several variations of the Borda count method in our upcoming sections, which we will refer to as BC variations. We must first distinguish between two implementations of the standard Borda count points system described by Emerson (2013) and used by McCune and McCune (2024) to account for partial or truncated ballots. The original points system introduced by Borda with the points vector $(n,n-1,\ldots,2,1)$ is assuming every ballot is complete; that is, each voter ranked a full list of candidates. In most ranked choice elections with $n$ candidates, however, voters may choose to submit a partial ballot in which they rank fewer than $n$ options. Additionally, some districts limit voters to ranking only a subset of candidates of a maximum size $m<n$. These are called truncated ballots. For example, the Minneapolis City Council elections only allow at most 3 candidates to be included in a ballot no matter how many candidates are running.
@@
The two primary approaches for handling partial or truncated ballots are the Borda count untreated (BCU) and the modified Borda count (MBC). If a ballot only ranks $k$ of the $n$ candidates, MBC awards $(k,k-1,\ldots,2,1,0,\ldots,0)$, so all unranked candidates receive 0 points. BCU awards points to the ranked candidates based on the full Borda count value as $(n,n-1,\ldots,n-k+1,0,\ldots,0)$. Emerson (2013) argues that BCU would encourage voters to use a form of tactical voting in which voters cast a partial ballot consisting of only their top choice, which is known as bullet voting. Giving a full $n$ points to that choice and 0 points to all other candidates would clearly give a voter’s top candidate the best chances of winning, and if enough voters follow this strategy, the election would devolve into a plurality vote and additionally lead to frequent truncation failures.
@@
One negative of both variations is that not all votes are considered equal. For example, in an election with $5$ candidates, a voter who fills out a complete ballot ranking all 5 candidates is assigning 15 total points to the candidates. A voter with a partial ballot of 3 candidates would only give candidates a total of 12 points using BCU and merely 6 points using MBC. While this does incentivize the electorate to fill out complete ballots, it is against democratic principles to count some voters more than others. Furthermore, in elections such as the 2021 East Hampton Mayor election, as many as 66 candidates may be present on the ballot. In elections with more than 8 candidates, which account for over 10 percent of the RCV elections in our dataset, it would not feasible to expect each voter to differentiate every candidate enough to provide a complete ranking without arbitrarily ordering the lower end of candidates. MBC in particular penalizes voters who want to submit a partial ballot to the point that they may be pressured to vote insincerely to ensure their top candidate is awarded the maximum number of points.
@@
To address this inequity, we favor the variation of the Borda count introduced by Dummett (1997) called the averaged Borda count (ABC). In the previous example with 5 candidates and a 3-candidate partial ballot, it initially follows BCU with 5 points for 1st, 4 for 2nd, and 3 for 3rd, but ABC will then average the unassigned 3 points to give 1.5 points to each of the unlisted candidates. In general, if $k$ candidates are ranked in an $n$-candidate election, they receive the same points as in BCU, while the unranked $n-k$ candidates receive an equal share of the remaining points that a complete ballot would have assigned. This equates to $\frac{1}{n-k}(1+2+\cdots+(n-k+1)+(n-k))=\frac{(n-k)(n-k+1)}{2(n-k)}=\frac{n-k+1}{2}$. In total, the points vector is
@@
The two points-scale variations that we introduce are the Exponential Borda Count (EBC) and Quadratic Borda Count (QBC). EBC awards 1 point to the last ranked candidate and double the points for each subsequently higher rank. This leads to the points for the $i$th candidate being the exponential function $p_{i}=2^{n-i}$, or points vector of $(2^{n-1},2^{n-2},\ldots,4,2,1)$. QBC also awards 1 point for a last-place vote and then the difference in points between ranks increases by 1 each time with the top two candidates having a difference of $p_{1}-p_{2}=n-1$. In general, $p_{i}-p_{i+1}=p_{i+1}-p_{i+2}+1$, and the formula for the points assigned to the $i$th ranked candidate is the quadratic function
@@
Table 2 shows the values of $p_{i}$ in a 6-candidate election to illustrate the differences in the variations.
@@
Now that we have defined the variations, we shift our focus to whether each voting failure can occur with each method. Table 3 displays this information for the currently used plurality and IRV methods, along with the five BC variations we are investigating. Any failure that is possibly exhibited by these variations will either occur within the empirical analysis in an upcoming section, or is demonstrated by the referenced example. The failures for which a Borda variation is immune will be proven so in the upcoming theorems, with the exception of MBC avoiding truncation failures, as this was previously shown (Kamwa 2022).
@@
The following are elections of very small voter size that demonstrate particular voting failures for some of our BC variations. Although these are toy examples to make the points calculation and pairwise comparisons easy, examples can be developed for much larger elections by keeping similar distributions in ballot types. As we will see in Section 6, none of these voting failure types would have occurred using the specified Borda variation on previous U.S. elections, hence the need for examples to demonstrate their possibility. We begin by demonstrating that EBC and QBC can exhibit a majority winner failure.
@@
Note the previous example is also a Condorcet winner failure, but we only focused on the majority aspect because it does not occur in our upcoming empirical analysis, whereas Condorcet winner failures are exhibited for every variation of the Borda count. The next example shows that EBC and QBC can both have majority loser and Condorcet loser failures. Although the Borda count cannot exhibit these two failures, as shown by Fishburn and Gehrlein (1976), any extension of Borda with a points vector where the difference $p_{i}-p_{i+1}$ increases as $i$ goes from $1$ to $n-1$ can have these failures if the number of voters is sufficiently large.
@@
Consider an election with candidates $A$, $B$, and $C$ using EBC or the equivalent QBC. Assume there are 4 ballots ranking $A>B>C$, 3 ranking $A>C>B$, 4 with $B>C>A$, and 4 with $C>B>A$. Notice that $A$ is ranked last in 8 of the 15 ballots, which is a majority and leads $A$ to lose in pairwise competition to both $B$ and $C$. Candidate $A$ wins the election, however, by earning 36 points while $B$ earns 35 and $C$ receives 34. This election exhibits both a majority loser and Condorcet loser failure.
@@
It was shown (Fishburn and Gehrlein 1976) that the Borda count, assuming all ballots are complete, cannot have a Condorcet loser failure. This implies a majority loser failure is also impossible. We now show two variations of Borda, particularly MBC and ABC, are also immune to majority loser failures, and in the case of ABC, Condorcet loser failures as well.
@@
We follow a similar approach to the previous result by supposing for a contradiction that there is an ABC election with a Condorcet loser failure. Assume $A$ is the winning candidate who loses in pairwise competition to all other candidates. The standard Borda count does not have Condorcet loser failures, so there must exist partial ballots.
@@
In the new Borda count election, all previously complete ballots are included twice. Any ballot where one of $A$ or $B$ was ranked and the other unranked would also count twice toward the pairwise comparison. This is because the originally ranked candidate is favored in both new ballots, despite the order of the newly ranked candidates being different in the two forms of completing ballots. This accounts for all cases where one of $A$ or $B$ was ranked higher, and thus there are so far $2x$ ballots with $B$ favored over $A$, and $2y$ with the opposite relative order. Assuming there were $z$ partial ballots that left $A$ and $B$ unranked, the two forms of ballot completion with reverse ordering result in $A$ ranked higher in $z$ newly completed ballots and $B$ also being higher in $z$ ballots. Now that we have covered all cases of both being ranked, one ranked and one unranked, and both unranked, $B$ is favored over $A$ in $2x+z$ ballots, while $A$ is favored in $2y+z$ ballots. The fact that $x>y$ directly implies $2x+z>2y+z$, meaning that $B$ wins the head-to-head comparison.
@@ 4 Data and Methodology
The election data used in our empirical analysis consists of 421 RCV elections from a publicly available database (Otis 2022). Table 4 summarizes the dataset, including the states and cities where the elections occurred and the range of years for each jurisdiction. The elections include local city council, school board, and mayor offices in addition to statewide elections for the U.S. House of Representatives and Presidential primaries.
In using the RCV data where voters chose their ranking based on IRV deciding the winner, we recognize a potential methodological flaw by which some voter preferences may differ if a points-based method was used. Strategic voting, such as truncating a ballot to remove your favored candidate’s top competitor, would nevertheless have little effect on potential verifiable failures. Meanwhile, the purpose of voter manipulation failures is to pinpoint the potential for this exact behavior and identify it as an unfair outcome, erasing any methodological concerns.
There are multiple considerations needed to process ballots and implement our points-based methods. For instance, there are many ballots in which a candidate was listed multiple times, ranks were skipped, or multiple candidates were listed in the same ranking, identified as an overvote. Any appearance of a candidate after the first instance was removed from the ballot, likewise for skipped or undeclared ranks. Then these were processed as partial ballots. Finally, for cases of overvotes, there was no way to determine which candidates were in the overvote, so all candidates below that ranking were removed, leaving only the candidates ranked above the overvote rank.
Another determination that must be made when implementing a points-based voting system is how to enumerate ballots that include write-in candidates. Across all elections in which write-in candidates were allowed to be included, they received no more than $1.6\%$ of the total first-place votes. Moreover, this number includes all write-in candidates collectively since data from elections offices usually reports these as a single candidate named Write-In or a similar designation. Therefore, we opted for removing all write-in candidates since they had a negligible effect on the election outcome.
@@
The presence of voter/candidate manipulation failures are more difficult to identify from a computational standpoint. For a truncation failure, our algorithm considers each losing candidate and truncates a set of ballots to give them an optimal chance of overtaking the winning candidate’s point total. This is accomplished by removing all ranks starting with the initial winner from each ballot where the particular losing candidate was ranked higher. If that candidate now wins, there is a truncation failure. There are scenarios, however, where a third candidate would now win, forcing us to further analyze those examples to determine if a smaller subset of ballots or if a different truncation strategy, such as removing this third candidate as well, would trigger a failure. Compromise failures follow a similar algorithmic process in which a losing candidate ranked higher than the winner is moved to the top of all such ballots to determine if they would now win.
To identify spoiler effects, we use a brute-force approach for elections with slates of 10 or fewer candidates. By sequentially removing each subset of losing candidates, the algorithm determines if the winner has changed, resulting in a spoiler effect. Unfortunately, it becomes computationally infeasible to perform this evaluation for every subset if there is a larger slate. To guarantee accuracy, we limited our analysis to the 396 elections with 10 or fewer candidates.